OpenStudy (anonymous):

A man bought 'a' eggs at 'r' cents per dozen. He sold them for 's' cents each. Find an expression, in terms of a,r and s, for the profit in cents, that he made

OpenStudy (anonymous):

@Callisto @satellite73 @ParthKohli @ash2326

OpenStudy (ash2326):

Let's find the price( Cost Price) he paid for the eggs \[CP=\frac{a}{12} \times r\] Do you get this?

OpenStudy (anonymous):

okay..

OpenStudy (ash2326):

He sold a eggs for s cents each, so selling price (SP) \[SP= a\times s\] Profit in cents= \[\text{Profit in Cents}=SP-CP=a\times s-\frac{a}{12}\times r\]

OpenStudy (anonymous):

can u simplify that?

OpenStudy (ash2326):

\[\text{Profit in Cents}=a\times s-\frac{a}{12}\times r=a \times (s-\frac{r}{12})\]

OpenStudy (anonymous):

\[\frac{a(12s-r)}{12}\---> this is the answer, how do we get this?

OpenStudy (anonymous):

\[\frac{a(12s-r)}{12}\] this one

OpenStudy (ash2326):

\[\text{Profit in Cents}=a\times s-\frac{a}{12}\times r=a \times (s-\frac{r}{12})=a(\frac{s\times 12}{1\times 12}-\frac{r}{12})\]\[=a(\frac{12s-r}{12})\]

OpenStudy (ash2326):

Do you understand this?

OpenStudy (anonymous):

oh okay u subtracted right and got that 12s- r , yeah i get it